OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of y = 2x – 4? A. 2x + y = 4 B. 4x – 2y = –12 C. y = –4x + 8 D. y = 3x – 4

OpenStudy (selrachcw95):

Srry have no clue but will guess c or b.

OpenStudy (anonymous):

Thanks help. Gave you a medal, will you fan me so I can message you?

OpenStudy (anonymous):

Yeah sure.

OpenStudy (anonymous):

Was d correct?

OpenStudy (anonymous):

Yes. I think so.

OpenStudy (shamil98):

You are wrong, parallel lines have the same slope ...

OpenStudy (shamil98):

The answer is b. 4x-2y=-12 -2y = -4x - 12 y = 2x + 6

OpenStudy (anonymous):

Thanks Shamil, I'd give you best answer but I gave to help already, but I got more questions as well

OpenStudy (anonymous):

Just click the UNDO BUTTON next to best response.

OpenStudy (anonymous):

And then give BEST RESPONSE to shamil!

OpenStudy (anonymous):

Sorry I could not make myself useful,Vinnie. Truth is,I'm terrible at it:(